### v1

# Injecting Domain-Specific Knowledge into Large Language Models: A Comprehensive Survey

Abstract
[…]
1 Introduction
Large Language Models (LLMs) have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making. However, despite their impressive performance in general-purpose applications, many specialized domains, such as healthcare, chemistry, and legal analysis, demand the integration of domain-specific knowledge to achieve high accuracy and reliability. To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection. This approach aims to bridge the gap between general-purpose language understanding and the stringent requirements of domain-specific tasks, enabling LLMs to perform effectively in highly specialized contexts.
Figure 1: Illustration of Growth Trends in Domain-Specific Knowledge Injection into LLMs. The chart displays the cumulative number of papers published between October 2022 and December 2024. Different colors and border styles represent various injection methods and domains, such as blue with a solid border denoting dynamic injection in the biomedical field.
Building on the foundational capabilities of general-purpose LLMs, knowledge injection techniques provide an effective means to address their limitations in handling specialized applications. Compared to the generalized approach of standard LLMs, knowledge injection offers two key advantages: 1) incorporating precise, domain-specific knowledge to improve accuracy and reliability in specialized tasks, and 2) allowing LLMs to dynamically adapt to new information or evolving knowledge bases, ensuring up-to-date expertise. These techniques bridge the gap between general-purpose understanding and domain-specific demands by leveraging both structured and unstructured knowledge sources. As a result, knowledge injection methods have been successfully applied in fields such as healthcare, chemistry, and legal analysis, significantly enhancing LLM performance. For example, biomedical LLMs  (Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024a) have achieved advances in material property prediction and discovery. These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs, unlocking solutions to complex, field-specific challenges.
Despite these advancements, early efforts in knowledge injection often treated domains independently, leading to a lack of standardization in methodologies and evaluation. As the volume of research continues to grow rapidly, with applications and studies proliferating across disciplines, the need for a comprehensive review becomes evident. This review aims to summarize the state of knowledge injection techniques, provide a systematic blueprint for future research, and identify key challenges, such as balancing scalability with domain-specific accuracy and enabling efficient, real-time knowledge updates.
To assist readers from various backgrounds in understanding knowledge injection techniques for LLMs and to complement existing surveys by addressing unresolved questions, we organize our survey paper as follows. After providing the foundational background knowledge in Section 2, we address a pivotal question: How can domain-specific knowledge be effectively integrated into LLMs? To answer this, we present a comprehensive framework for categorizing knowledge injection methods in Section 3. We delve into this topic by discussing: 1) Dynamic Knowledge Injection, which explains how external knowledge is retrieved and incorporated in real-time during inference to enhance reasoning capabilities; 2) Static Knowledge Embedding, which describes how domain knowledge is embedded into the model during training or fine-tuning to make it an inherent part of the model; 3) Adapters, which highlight modular techniques for storing and utilizing external knowledge without altering the primary model’s parameters; and 4) Prompt Optimization, which focuses on how carefully designed prompts enable models to utilize existing knowledge without requiring changes to their architecture. Another perspective for reviewing knowledge injection studies is their application across various domains. In Section 4, we categorize current application domains, such as materials science, chemistry, biology, and law. To guide the identification of appropriate tools and resources, we summarize commonly used benchmarks, open-source frameworks, and analyses on performance in Section 5. Based on these summaries, we discuss the challenges and opportunities in this evolving field in Section 6. The conclusions are summarized in Section 7.
2 Background
2.1 Domain-Specific Knowledge
Domain-specific knowledge refers to specialized information or expertise pertinent to a specific field or application, distinguishing it from general knowledge that spans across multiple domains. While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required. For instance, in scientific text processing (Bran et al., 2023), models must comprehend complex scientific terminologies, concepts, and methodologies to provide accurate and relevant answers. Similarly, in e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations. In healthcare applications, LLMs must understand medical terminologies, diagnoses, treatment plans, and drug interactions. For example, biomedical question answering (Pei and others, 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017).
To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs. In this paper, we aim to provide a survey of these various injection methods.
2.2 Knowledge Representation and Encoding
Knowledge can take different forms depending on the structure and application needs. For example, knowledge graphs (Zhang et al., 2024c) represent information as entities and relationships in a graph format, allowing structured reasoning and inference. These are widely used in tasks like question-answering and recommendation systems, where relationships between entities are crucial. Similarly, knowledge in text form, such as Wikipedia (Jeong et al., 2024), provides a vast corpus of unstructured information.
Knowledge can also be stored in vector space instead of readable text or graph formats. For instance, soft prompt tuning (Singhal et al., 2023a) learns useful knowledge in vector form, which is concatenated with the original input to guide LLMs in performing specific downstream tasks. In addition to external representations, knowledge can also emerge from within the model itself. For example, chain-of-thought prompting (Yao et al., 2024) introduces intermediate reasoning steps that help the model break down complex tasks into manageable parts. By explicitly reasoning through these steps, the LLM can utilize its internally stored information more effectively, resulting in better performance on tasks requiring logical reasoning, multi-step computation, or decision-making.
3 Paradigms of Knowledge Injection
This section identifies four major paradigms of knowledge injection: Dynamic Knowledge Injection, Static Knowledge Embedding, Adapters, and Prompt Optimization. These paradigms illustrate various mechanisms by which external domain-specific knowledge can be incorporated into LLMs. We utilize unified notations, as described in Table 1, to systematically represent the processes. External knowledge $\mathcal{K}$ is integrated into LLMs either by modifying the original parameters $\theta$, introducing additional parameters $\phi$, or leveraging auxiliary mechanisms.
Figure 2: Four knowledge injection paradigms for LLMs. (a) Dynamic Knowledge Injection retrieves external knowledge during inference for enhanced reasoning. (b) Static Knowledge Injection embeds external knowledge into model parameters during fine-tuning. (c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks or updates. (d) Prompt Optimization utilizes precise prompts to guide the LLM without altering its parameters.
3.1 Dynamic Knowledge Injection
Symbol Description
$\mathbf{x}$ Input to LLM
$\mathbf{y}$ Output of LLM
$M$ Backbone LLM Function
$\mathcal{K}$ External domain knowledge base
$\theta$ Parameters of LLM
$\phi$ Additional parameters introduced
$\mathcal{R}(\mathbf{x},\mathcal{K})$ Retrieval function fetches relevant elements of $\mathcal{K}$ given the input $\mathbf{x}$
$M(\mathbf{x};\theta)$ Represent LLM takes input $\mathbf{x}$ and produces an output, parameterized by $\theta$
$\Delta\theta$ Offsets to the original LLM’s parameters
Table 1: Summary of Symbols.
We define dynamic knowledge injection as the process of first retrieving information from external knowledge bases or knowledge graphs and then combining it with the input for use in LLMs:
$\mathbf{y}=M(\mathbf{x},\mathbf{\mathcal{R}(\mathbf{x},\mathcal{K})};\theta),$ (1)
where $\mathbf{x}$ represents the original input, $\mathcal{R}$ denotes the retrieval function, $\mathcal{K}$ is the external knowledge base, and $\theta$ are the model parameters, which remain unchanged. This paradigm offers several advantages, including ease of updating (hence the term ”dynamic injection”) and the ability to incorporate new knowledge without retraining the model. However, it also presents challenges, such as dependency on the quality of the knowledge base $\mathcal{K}$, the retrieval function $\mathcal{R}$, and limitations imposed by the maximum input length of the LLM.
3.2 Static Knowledge Embedding
Compared with dynamic knowledge retrieval, static knowledge embedding involves embedding knowledge into the model’s parameters through full or partial fine-tuning, making it less flexible to changes. Concretely, the model learns new parameters $\Delta\theta$ that encode domain knowledge from $\mathcal{K}$:
$\Delta\theta=\textstyle\arg\min_{\theta}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta),\mathbf{y_{s}}\bigr{)},$
where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a typical supervised training loss function. After optimization, the updated parameters $\Delta\theta$ are obtained.
At inference time, no further retrieval or external knowledge calls are required:
$\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}.$
This paradigm provides fast inference, as it eliminates additional retrieval steps and often results in stronger performance. However, it also comes with challenges, such as costly updates—requiring fine-tuning whenever domain knowledge changes—and scalability issues, as embedding large or frequently changing knowledge bases can be computationally expensive.
3.3 Modular Knowledge Adapters
To address the costly updates associated with static knowledge embedding, another paradigm, known as modular knowledge adapters, introduces small, trainable modules that can be inserted into or operate alongside the base model to store domain-specific knowledge while saving computational resources. In this approach, the original parameters $\theta$ of the LLM typically remain frozen, preserving the model’s general-purpose capabilities. Given a knowledge dataset $\mathcal{K}$, the adapter parameters $\phi$ are trained by minimizing the following objective:
$\phi=\textstyle\arg\min_{\phi}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta,\phi),\mathbf{y_{s}}\bigr{)},$
where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters. At inference time, the enhanced model generates outputs as:
$\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}.$
This paradigm offers a parameter-efficient method to adapt LLMs to specific domains without modifying the original model weights. By freezing the base model’s parameters, the approach seeks to preserve previously acquired knowledge while enabling the seamless incorporation of new domain-specific information. However, this approach also introduces challenges, such as the need to design new architectural components and determine appropriate hyperparameters, including the size and number of adapters. These additional elements can increase the overall complexity of the model and its training process.
3.4 Prompt Optimization
Unlike previous approaches, prompt optimization does not retrieve knowledge from external sources. Instead, it focuses on fully leveraging or guiding the LLM to utilize its internal, pre-existing knowledge. The process can be formalized as:
$\mathbf{y}=M\bigl{(}[\mathbf{p},\mathbf{x}];\theta\bigr{)},$
where $\mathbf{p}$ represents a textual prompt containing implicit domain knowledge or specific instructions.
Prompt optimization offers significant advantages, including eliminating dependency on external domain knowledge bases and avoiding training. However, it also presents challenges, as designing effective prompts can be both complex and time-consuming. Additionally, long prompts may reduce the available context window, potentially affecting model efficiency and performance.
3.5 Comparison of the Four Paradigms
Paradigm Training Cost Inference Speed Limitations
Dynamic
Injection
None, but requires
retrieval module
Slower due to
retrieval latency
Relies heavily on
retrieval quality
Static
Embedding
High
(requires pretraining
or fine-tuning)
No extra cost
Fixed knowledge;
risks catastrophic
forgetting
Modular
Adapters
Low
(train small subset
of parameters)
Almost
unaffected
Sensitive to training
data quality
Prompt
Optimization
None
Almost
unaffected
Labor-intensive;
limited to pre-existing
knowledge
Table 2: Guidance on selecting knowledge injection paradigms based on training cost, inference speed, and limitations.
Dynamic knowledge injection integrates external knowledge at runtime, offering flexibility and adaptability to new information without increasing training cost. However, it requires an effective retrieval module, and the inference speed depends heavily on retrieval performance, which can slow down the overall process. Static knowledge embedding embeds domain expertise during pretraining or fine-tuning, requiring large-scale domain-specific data and significant training resources, including GPUs and time. While it incurs no additional inference cost, its limitations lie in the potential risks of catastrophic forgetting and its inability to adapt to evolving information. Modular adapters serve as a middle ground, allowing plug-and-play components to enhance domain-specific capabilities with minimal training data. Only a small subset of parameters needs to be trained, which reduces training costs, and inference speed is largely unaffected. However, the quality of training data significantly impacts the performance of this method. Prompt Optimization, on the other hand, avoids retraining entirely by leveraging carefully crafted inputs. It has no impact on inference speed but relies on extensive human effort to find optimal prompts. This method is limited in its ability to utilize new knowledge and primarily activates pre-existing knowledge.
We summarize these comparisons in Table 2 as a practical guide to help determine the most suitable method based on specific task requirements and scenarios.
4 Applications
4.1 Biomedicine
The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts. These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data. For instance, PMC-LLaMA (Wu et al., 2023) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively. Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning. This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
Beyond foundational models, integrating external tools and knowledge can further enhance performance. For example, GeneGPT (Jin et al., 2024) utilizes an LLM pretrained on code tasks to tackle GeneTuring tests by employing NCBI Web APIs. This approach combines in-context learning with an augmented decoding algorithm that identifies and executes API calls. Similarly, Med-PaLM (Singhal et al., 2023a) introduces vector prompts—representations that store and retrieve medical domain knowledge—to extend the capabilities of Flan-PaLM (Chung et al., 2024).
Domain Model Paradigms Knowledge Source Link
Biomedicine PMC-LLaMA (Wu et al., 2023) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
Med-PaLM 2 (Singhal et al., 2023b) Static Knowledge Embedding MultiMed Link
DALK  (Li and others, 2024)
Dynamic Knowledge Injection
Prompt Optimization
MedQA, MedMCQA, MMLU, QA4MRE Link
ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
SA-MDKIF (Xu et al., 2024b) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
MaLP (Zhang et al., 2024b) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
BioMedLM  (Bolton et al., 2024) Static Knowledge Embedding PubMed,MedMCQA,MedQA,MMLU,BioASQ Link
BiomedRAG (Li et al., 2024) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
Finance FLANG  (Shah and others, 2022) Static Knowledge Embedding
Financial PhraseBank,FiQA 2018 Task-1,
News Headline Classification, Named Entity Recognition,
Structure Boundary Detection,Question Answering
Link
BloomBergGPT (Wu and others, 2023)
Static Knowledge Embedding
Finance dataset (web, news, filings, press, Bloomberg),
Public dataset (the Pile, C4, Wikipedia)
FinMA  (Xie et al., 2023) Static Knowledge Embedding
FPB,FiQA-SA,Headline,NER,FinQA,
ConvFinQA,BigData22,ACL18,CIKM18
Link
FinGPT (Zhang et al., 2023) Modular Knowledge Adapters
Financial news, Company filings and announcements,
Social media discussions, Trends
Link
Fin-LLaMA  (Konstantinidis and others, 2024) Static Knowledge Embedding fin-llama-dataset Link
SNFinLLM (Zhao et al., 2024a)
Static Knowledge Embedding
FinEval, FinanceIQ,qEQA,FinC,KQA,MRC,cMRC
Materials ChemCrow (Bran et al., 2023) Dynamic Knowledge Injection 18 expert-designed tools Link
ChemDFM  (Zhao et al., 2024b) Static Knowledge Embedding
SciQ,PIQA,PubChem,ARC,USPTO
Link
ChemLLM (Zhang and others, 2024)
Static Knowledge Embedding
ChemData,ChemBench
Link
CrystaLLM  (Antunes et al., 2024)
Static Knowledge Embedding
Materials Project, OQMD, NOMAD
Link
ScholarChemQA  (Chen and others, 2024)
Static Knowledge Embedding
AG News,Yahoo Answers ,Yelp-5,Amazon-5 Link
DARWIN 1.5 (Xie et al., 2024) Static Knowledge Embedding FAIR datasets Link
ChemAgent (Tang and others, 2025)
Dynamic Knowledge Injection
Prompt Optimization
Google API, Docker container, Internet,
Hardware API document, Physical world hardware
Link
Mental Health MeChat (Qiu et al., 2023) Dynamic Knowledge Injection SMILECHAT, PsyQA Link
MindChat (Xin Yan, 2023) Static Knowledge Embedding Multi-turn psychological dialogue data Link
SoulChat (Chen et al., 2023) Static Knowledge Embedding Long-text counseling sessions Link
EmoLLM (Yang et al., 2024)
Static Knowledge Embedding
Modular Knowledge Adapters
CPsyCounD Link
Education EduChat (Dan et al., 2023) Static Knowledge Embedding
Textbooks Data, Open QA Data,
Emotional Support Data, Socratic Teaching Data
Link
QiaoBan (Weixiang et al., 2023) Prompt Optimization Children’s emotional education dialogue data Link
HiTA (Liu et al., 2024a) Dynamic Knowledge Injection Educator curated database \
SocraticLM (Liu et al., 2024c) Modular Knowledge Adapters SocraTeach dataset \
CyberQ (Agrawal et al., 2024)
Static Knowledge Embedding
Dynamic Knowledge Injection
AISecKG, Q&A \
Social Science SocialLLM (Jiang and Ferrara, 2023)
Static Knowledge Embedding
Prompt Optimization
Covid-Political, Election2020, COVID-Morality,
Ukr-Rus-Suspended, Ukr-Rus-Hate,
Immigration-Hate-08, Immigration-Hate-05
FPS (Liu et al., 2024e) Prompt Optimization Fake News Dataset, Big Five Personality Traits Link
FUSE (Liu et al., 2024f) Prompt Optimization True News Dataset, Big Five Personality Traits \
Table 3: Summary of the domain-specific knowledge injection studies. We categorize current work according to their research domain and knowledge injection method.
4.2 Finance
Fine-tuned financial LLMs have demonstrated significant advancements in adapting general-purpose models to domain-specific tasks through task-specific training. PIXIU (Xie et al., 2023) fine-tunes LLaMA on 136K instruction samples tailored to financial tasks, enabling the model to handle a wide range of domain-relevant scenarios. Instruct-FinGPT (Zhang et al., 2023) fine-tunes LLaMA on 10K instruction samples derived from two financial sentiment analysis datasets, focusing primarily on finance classification tasks. FinGPT (Yang et al., 2023) introduces a comprehensive end-to-end framework for training and deploying FinLLMs in the financial industry. Utilizing the LoRA technique, FinGPT fine-tunes open-source LLMs like LLaMA and ChatGLM with approximately 50K task-specific samples, achieving efficient fine-tuning without full model retraining.
In contrast, scratch-trained financial LLMs aim to create models specifically designed for financial tasks from the ground up. BloombergGPT (Wu and others, 2023) leverages a subset of 5 billion tokens from Bloomberg-specific data, representing only 0.7% of its total training corpus, to tailor its model to financial applications. XuanYuan 2.0 (Zhang and Yang, 2023) combines 366 billion tokens for pre-training with an additional 13 billion tokens for fine-tuning, creating the largest Chinese financial chat model. Similarly, Fin-T5 (Lu et al., 2023) introduces a Chinese financial pre-training language model built on the T5 architecture, utilizing a 300GB financial corpus. Furthermore, SNFinLLM (Zhao et al., 2024a) dynamically incorporates real-time financial data during inference to enhance decision-making capabilities, demonstrating the value of domain-specific pretraining and adaptability in financial LLMs.
4.3 Materials
In contrast to biomedicine, where significant efforts have been devoted to static knowledge embedding due to the availability of extensive corpora, research in materials and chemistry has primarily focused on utilizing task-related tools that align with the dynamic knowledge injection paradigm.
For instance,  Xie et al. (2024) demonstrated how Darwin 1.5 leverages natural language inputs and a two-stage training strategy to achieve significant improvements in materials discovery and design tasks. Bran et al. (2023) introduced ChemCrow, a framework that augments LLMs with chemistry-expert-designed tools for downstream tasks such as organic synthesis and drug discovery. There are also studies on prompt optimization Tang and others (2025), demonstrating that better-designed planning prompts can effectively utilize the model’s internal knowledge to orchestrate complex tasks. This approach capitalizes on the planning and execution capabilities of multiple LLMs to achieve autonomy in chemical experimentation.
More recently, there has been growing interest in exploring static knowledge embedding and modular knowledge adapters within the chemistry domain. For example, Chen and others (2024) curated a QA dataset to fine-tune pretrained models like BERT and LLMs such as Llama, aiming to enhance their performance in chemistry-related tasks. Similarly, Xie et al. (2024) introduced Darwin 1.5, an open-source large language model tailored for materials science.
4.4 Human-Centered Science
The last domain we introduce is human-centered science, which encompasses a wide range of applications such as psychological counseling, financial forecasting, social behavior prediction, and legal reasoning. All these domains center on understanding and interacting with human needs, behaviors, and decision-making processes.
In mental health, datasets like PsyQA (Sun et al., 2021) provide a foundation for training models in psychological counseling scenarios. For instance, SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations. Similarly, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, significantly enhancing its emotional support capabilities. These advancements demonstrate the potential of human-centered science in addressing complex, real-world challenges through personalized and context-aware solutions.
In the education domain, LLMs have shown immense potential in addressing challenges such as personalized learning, curriculum alignment, and interactive teaching. Personalized learning, for example, requires models to adapt to individual needs, providing tailored feedback and emotional support. EduChat (Dan et al., 2023) tackles this by leveraging educational theories from psychology and pedagogy through static knowledge embedding, enabling tasks like open Q&A, composition correction, and emotional support. Similarly, QiaoBan (Weixiang et al., 2023) focuses on child-centered education by using prompt optimization to adapt the model’s behavior based on child psychology and emotional well-being, catering specifically to young learners. Domain-specific education and interactive teaching have also seen advancements through LLMs. CyberQ (Agrawal et al., 2024) blends static knowledge embedding and dynamic knowledge injection via AISecKG (Agrawal, 2023), generating Q&A based on cybersecurity best practices. Interactive teaching, on the other hand, benefits from models like SocraticLM (Liu et al., 2024c), which employs adapters fine-tuned on the SocraTeach dataset to engage students in critical thinking and problem-solving.
For social sciences, models like SocialLLM (Jiang and Ferrara, 2023) combine static knowledge embedding and dynamic knowledge injection to analyze human behavior in social networks. Adapters facilitate large-scale data integration while prompt optimization guides the model to focus on specific social behavior patterns. Models like FPS (Liu et al., 2024e) and FUSE (Liu et al., 2024f) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact.
A summary of the mainstream models and their information is provided in Table 3. More models across various domains can be found at: Survey-official-repo.
5 Tools, Resources, and Analysis
5.1 Knowledge Injection Framework
In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024c) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
KnowGPT dynamically combines knowledge graphs with prompt optimization by leveraging reinforcement learning to extract highly relevant subgraphs from the knowledge graph. These subgraphs are represented as triples and transformed into natural language prompts that language models can interpret and utilize via diverse prompt templates. The KnowGPT framework significantly reduces the API call costs of LLMs while enhancing their performance in domain-specific tasks.
StructTuning uses a structure-aware approach to embed domain knowledge into pre-trained models with a two-stage strategy: Structure-Aware Continual Pre-Training encodes knowledge into the model’s parameters, and Structure-Aware Supervised Fine-Tuning refines understanding through structured QA tasks. This framework demonstrates significant performance improvements in knowledge-driven tasks such as relation classification and question answering, achieving a balance between generality and efficiency.
K-Adapter stores knowledge within adapter modules. Its core method involves freezing the original model parameters and assigning an independent, task-specific adapter for each type of knowledge. These adapters are inserted as independent modules into the intermediate layers of the model to generate enhanced representations of specific knowledge. This design effectively mitigates the issue of catastrophic forgetting, preventing newly injected knowledge from overwriting the model’s pre-existing knowledge.
Finally, SelfLift iteratively employs a retrieval-augmented generator to create an unbounded memory pool and uses a memory selector to choose one output as memory for the subsequent generation round. This is an excellent demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
5.2 Datasets and Benchmarks
We summarize commonly used datasets or benchmarks for domain-specific LLM study in Table 3, observing significant variation in dataset richness across domains. Biomedicine boasts numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis and others (2012), that support tasks such as question answering and clinical summarization. In contrast, materials and chemistry have more limited resources, and datasets like USPTO and Enzymes focus on chemical reactions. Miscellaneous datasets are scattered across other domains, such as PsyQA and SmileChat in mental health, SocraTeach, and Children’s emotional education dialogue data dataset in education. This diversity underscores the effort to tailor LLMs to specialized fields while emphasizing the need for broader curation of benchmarks in underrepresented domains.
5.3 Performance Comparison of Domain-specific LLM and General-domain LLM
Model Domain Type Size MedQA PubMedQA MedMCQA
Med-Gemini Specific 91.1 - -
GPT-4 General 90.2 80.4 73.7
Med-PaLM 2 Specific 85.4 81.8 72.3
PMC-LLaMA Specific 13B 56.3 77.9 56.0
BioMedLM Specific 2.7B 50.3 74.4
Llama 2 General 70B 43.7 74.3 35.0
Galactica General 120B 44.4 77.6 52.9
Table 4: Performance comparison of domain-specific and general-domain model performance on medical benchmarks.
Since there are also general-domain strong LLMs, it is necessary to discuss the comparison between domain-specific LLMs and general ones to determine if a specific knowledge injection process is essential. Here, we take the biomedical domain as an example due to significant research efforts in this field, as shown in Table 4. The results are collected from corresponding papers or paperswithcode.com.
First, we can observe that closed-source LLMs are currently the most effective models, while the performance gap between general-domain and domain-specific LLMs is relatively narrow. For example, both GPT-4 and Med-Gemini (Saab et al., 2024) achieve outstanding performance, with scores higher than 90 on the MedQA dataset. However, due to the lack of transparency in closed-source LLMs, open-source LLM efforts should not be overlooked. In this field, domain-specific LLMs often outperform general-domain models. For instance, PMC LLaMA-13B outperforms LLaMA2-70B by more than 10 points on the MedQA dataset. This demonstrates the value of domain-specific LLMs in achieving superior performance on specialized tasks. While general-domain models can deliver strong results, incorporating domain-specific knowledge allows for significant improvements, particularly in open-source initiatives. This underscores the importance of investing in domain-specific LLMs to address the unique challenges of specialized fields.
6 Challenges and Opportunities
6.1 Integrated Knowledge Consistency
Knowledge injection allows LLMs to incorporate and integrate different domain-specific knowledge. However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024a). For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise, resulting in unreliable decisions and undermining the system’s trustworthiness.
To address this, future research must focus on detecting inconsistencies, resolving conflicts, and maintaining consistency in integrated knowledge. Conflicts can be addressed by prioritizing reliable sources, applying domain-specific rules, or using ensemble techniques to balance multiple perspectives. Alignment algorithms and validation modules can further ensure that retrieved knowledge aligns with the model’s reasoning processes and is reliable before influencing outputs. These efforts are essential to enhance the reliability and applicability of knowledge-enhanced LLMs in complex, high-stakes domains.
6.2 Cross-Domain Knowledge Transfer
Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields. While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns. For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks.
Overcoming these challenges necessitates advancements in modular knowledge representation and transfer learning techniques. Future efforts could explore hybrid approaches that blend static embeddings with dynamic retrieval, enabling LLMs to adapt knowledge flexibly across domains without compromising depth. Additionally, creating standardized cross-domain benchmarks and datasets could facilitate systematic evaluation and foster innovation in multi-domain knowledge transfer methodologies.
7 Conclusion
LLMs enhanced by domain-specific knowledge have shown remarkable potential and garnered increasing research interest. This survey systematically reviews LLM knowledge injection systems, exploring knowledge representation methods, integration strategies, and mechanisms for preserving model generality. We also summarize applications across biomedicine, chemistry, and computational social science domains. By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires further exploration of knowledge-enhanced LLMs for domain-specific challenges.

### v2

@@
 […]
+Injecting Domain-Specific Knowledge into Large Language Models:
+A Comprehensive Survey
+Zirui Song1∗   Bin Yan2thanks: Equal contributions.
+Yuhan Liu3   Miao Fang2   Mingzhe Li4   Rui Yan5   Xiuying Chen1thanks: Corresponding author.
+1Mohamed bin Zayed University of Artificial Intelligence       2Northeastern University       3Gaoling School of Artificial Intelligence, Renmin University of China       4ByteDance       5Wuhan University E-mail: zirui.song@mbzuai.ac.ae, xiuying.chen@mbzuai.ac.ae
 1 Introduction
-Large Language Models (LLMs) have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making.
+LLMs have achieved extraordinary success across various tasks, showcasing remarkable capabilities in reasoning, knowledge representation, and decision-making Song et al.
+(2025); Xu et al.
+(2025).
 However, despite their impressive performance in general-purpose applications, many specialized domains, such as healthcare, chemistry, and legal analysis, demand the integration of domain-specific knowledge to achieve high accuracy and reliability.
-To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection.
+To address this challenge, researchers have explored methods to enhance LLMs through external or embedded domain expertise, a process often referred to as knowledge injection, as shown in Figure 1.
 This approach aims to bridge the gap between general-purpose language understanding and the stringent requirements of domain-specific tasks, enabling LLMs to perform effectively in highly specialized contexts.
@@
 The chart displays the cumulative number of papers published between October 2022 and December 2024.
-Different colors and border styles represent various injection methods and domains, such as blue with a solid border denoting dynamic injection in the biomedical field.
+Different colors and border styles represent various injection methods and domains.
 Building on the foundational capabilities of general-purpose LLMs, knowledge injection techniques provide an effective means to address their limitations in handling specialized applications.
@@
 As a result, knowledge injection methods have been successfully applied in fields such as healthcare, chemistry, and legal analysis, significantly enhancing LLM performance.
-For example, biomedical LLMs  (Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024a) have achieved advances in material property prediction and discovery.
-These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs, unlocking solutions to complex, field-specific challenges.
+For example, biomedical LLMs  (Cho and Lee, 2025; Bolton et al., 2024; Yan et al., 2023) have demonstrated superior accuracy in tasks like medical diagnostics and regulatory compliance, while domain-specific models for material science (Tang et al., 2025a; Xie et al., 2024; Antunes et al., 2024; Zhang et al., 2024b) have achieved advances in material property prediction and discovery.
+These dedicated models underscore the transformative potential of integrating domain knowledge into LLMs.
 Despite these advancements, early efforts in knowledge injection often treated domains independently, leading to a lack of standardization in methodologies and evaluation.
@@
 This review aims to summarize the state of knowledge injection techniques, provide a systematic blueprint for future research, and identify key challenges, such as balancing scalability with domain-specific accuracy and enabling efficient, real-time knowledge updates.
-To assist readers from various backgrounds in understanding knowledge injection techniques for LLMs and to complement existing surveys by addressing unresolved questions, we organize our survey paper as follows.
-After providing the foundational background knowledge in Section 2, we address a pivotal question: How can domain-specific knowledge be effectively integrated into LLMs?
-To answer this, we present a comprehensive framework for categorizing knowledge injection methods in Section 3.
-We delve into this topic by discussing: 1) Dynamic Knowledge Injection, which explains how external knowledge is retrieved and incorporated in real-time during inference to enhance reasoning capabilities; 2) Static Knowledge Embedding, which describes how domain knowledge is embedded into the model during training or fine-tuning to make it an inherent part of the model; 3) Adapters, which highlight modular techniques for storing and utilizing external knowledge without altering the primary model’s parameters; and 4) Prompt Optimization, which focuses on how carefully designed prompts enable models to utilize existing knowledge without requiring changes to their architecture.
-Another perspective for reviewing knowledge injection studies is their application across various domains.
-In Section 4, we categorize current application domains, such as materials science, chemistry, biology, and law.
-To guide the identification of appropriate tools and resources, we summarize commonly used benchmarks, open-source frameworks, and analyses on performance in Section 5.
-Based on these summaries, we discuss the challenges and opportunities in this evolving field in Section 6.
-The conclusions are summarized in Section 7.
+We begin in Section 2 with background on domain-specific knowledge and its role in LLMs.
+Section 3 presents a unified framework of four knowledge injection paradigms: (1) Dynamic Knowledge Injection at inference time; (2) Static Knowledge Embedding during training or fine-tuning; (3) Modular Adapters for parameter-efficient integration; and (4) Prompt Optimization via carefully designed inputs.
+Section 4 examines these methods across domains such as materials science, chemistry, biology, and law.
+Section 5 summarizes key datasets, tools, and comparative results.
+Section B outlines open challenges, including scalability, robustness, and domain transfer.
+Finally, Section 7 concludes the paper and reflects on future directions.
 2 Background
@@
 Domain-specific knowledge refers to specialized information or expertise pertinent to a specific field or application, distinguishing it from general knowledge that spans across multiple domains.
-While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required.
+While general knowledge enables models to understand broad contexts, domain-specific knowledge is essential for addressing specialized tasks where precise, field-specific understanding is required Wang et al.
+(2025); Li et al.
+(2025a); Zhang et al.
+(2025).
 For instance, in scientific text processing (Bran et al., 2023), models must comprehend complex scientific terminologies, concepts, and methodologies to provide accurate and relevant answers.
-Similarly, in e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations.
+In e-commerce search (Zhao et al., 2024a), understanding domain-specific terms such as product categories, technical specifications, or colloquial shopping language is crucial for delivering relevant search results and recommendations.
 In healthcare applications, LLMs must understand medical terminologies, diagnoses, treatment plans, and drug interactions.
-For example, biomedical question answering (Pei and others, 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017).
-To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs. In this paper, we aim to provide a survey of these various injection methods.
+For example, biomedical question answering (Singhal et al., 2025; Pei et al., 2024) and medical report summarization rely on integrating knowledge from medical literature like PubMed (Dernoncourt and Lee, 2017).
+To address these needs, researchers have explored various methods for incorporating domain-specific knowledge into LLMs.
+In this paper, we aim to provide a survey of these various injection methods.
 2.2 Knowledge Representation and Encoding
-Knowledge can take different forms depending on the structure and application needs.
-For example, knowledge graphs (Zhang et al., 2024c) represent information as entities and relationships in a graph format, allowing structured reasoning and inference.
-These are widely used in tasks like question-answering and recommendation systems, where relationships between entities are crucial.
-Similarly, knowledge in text form, such as Wikipedia (Jeong et al., 2024), provides a vast corpus of unstructured information.
-Knowledge can also be stored in vector space instead of readable text or graph formats.
-For instance, soft prompt tuning (Singhal et al., 2023a) learns useful knowledge in vector form, which is concatenated with the original input to guide LLMs in performing specific downstream tasks.
-In addition to external representations, knowledge can also emerge from within the model itself.
-For example, chain-of-thought prompting (Yao et al., 2024) introduces intermediate reasoning steps that help the model break down complex tasks into manageable parts.
-By explicitly reasoning through these steps, the LLM can utilize its internally stored information more effectively, resulting in better performance on tasks requiring logical reasoning, multi-step computation, or decision-making.
+Knowledge can take different forms depending on structure and application needs.
+Knowledge graphs (Liao et al., 2025; Zhang et al., 2024d) encode entities and their relationships in a structured graph, supporting reasoning and inference in tasks like question answering.
+In contrast, text-based sources like Wikipedia (Jeong et al., 2024) provide rich but unstructured information, useful for tasks requiring broad contextual understanding.
+Knowledge can also be stored in vector space rather than in text or graph form.
+For example, soft prompt tuning (Peng et al., 2025; Singhal et al., 2023a) embeds useful information as vectors, which are appended to inputs to guide LLMs on specific tasks.
+Beyond external forms, knowledge may also emerge internally: chain-of-thought prompting (Sanwal, 2025; Yao et al., 2024) introduces intermediate reasoning steps that help LLMs decompose complex problems and access internal knowledge more effectively—improving performance in tasks involving reasoning, multi-step computation, or decision-making.
+2.3 Knowledge Injection Survey
+Prior surveys on knowledge-enhanced language models vary in focus and scope.
+The most relevant works include the following: Cadeddu et al.
+(2024), who focus on scientific article classification and offer practical insights but lack broader methodological generalization; Wang et al.
+(2024b), who focus on knowledge editing and aim to update internal model knowledge with minimal side effects; and Hu et al.
+(2023), who adopt a model-centric perspective by classifying knowledge-enhanced models based on task type and knowledge source, though they primarily cover pre-LLM architectures such as BERT and ERNIE.
+In contrast, our work presents a unified view of knowledge injection in LLMs, emphasizing capability enhancement through external knowledge integration across diverse tasks.
 3 Paradigms of Knowledge Injection
-This section identifies four major paradigms of knowledge injection: Dynamic Knowledge Injection, Static Knowledge Embedding, Adapters, and Prompt Optimization.
-These paradigms illustrate various mechanisms by which external domain-specific knowledge can be incorporated into LLMs.
+To systematically understand how domain knowledge is integrated into LLMs, we categorize existing approaches into four paradigms based on when the knowledge is incorporated and how it interacts with the model, as shown in Figure 2.
+Specifically, Static Knowledge Injection and Modular Knowledge Adapters integrate knowledge prior to inference and involve parameter updates—through either full fine-tuning or adapter-based tuning.
+In contrast, Dynamic Knowledge Injection and Prompt Optimization inject knowledge at inference time without altering model parameters: the former retrieves external information, while the latter leverages internal knowledge through designed prompts.
 We utilize unified notations, as described in Table 1, to systematically represent the processes.
-External knowledge $\mathcal{K}$ is integrated into LLMs either by modifying the original parameters $\theta$, introducing additional parameters $\phi$, or leveraging auxiliary mechanisms.
 Figure 2: Four knowledge injection paradigms for LLMs.
-(a) Dynamic Knowledge Injection retrieves external knowledge during inference for enhanced reasoning.
+(a) Dynamic Knowledge Injection retrieves external knowledge during inference.
 (b) Static Knowledge Injection embeds external knowledge into model parameters during fine-tuning.
-(c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks or updates.
+(c) Modular Knowledge Adapters use plug-and-play modules to dynamically adapt to tasks.
 (d) Prompt Optimization utilizes precise prompts to guide the LLM without altering its parameters.
@@
 where $\mathbf{x}$ represents the original input, $\mathcal{R}$ denotes the retrieval function, $\mathcal{K}$ is the external knowledge base, and $\theta$ are the model parameters, which remain unchanged.
-This paradigm offers several advantages, including ease of updating (hence the term ”dynamic injection”) and the ability to incorporate new knowledge without retraining the model.
+This paradigm offers several advantages, including ease of updating (hence the term "dynamic injection") and the ability to incorporate new knowledge without retraining the model.
 However, it also presents challenges, such as dependency on the quality of the knowledge base $\mathcal{K}$, the retrieval function $\mathcal{R}$, and limitations imposed by the maximum input length of the LLM.
+To improve retrieval quality, commonly used techniques include semantic matching based on sentence embeddings and efficient vector databases for fast similarity search.
 3.2 Static Knowledge Embedding
@@
 $\Delta\theta=\textstyle\arg\min_{\theta}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta),\mathbf{y_{s}}\bigr{)},$
-where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a typical supervised training loss function.
+where $\mathcal{K}$ is the domain-specific knowledge base containing training samples $\mathbf{x_{s}}$ and $\mathbf{y_{s}}$, and $\mathcal{L}$ is a training loss function.
 After optimization, the updated parameters $\Delta\theta$ are obtained.
-At inference time, no further retrieval or external knowledge calls are required:
-$\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}.$
-This paradigm provides fast inference, as it eliminates additional retrieval steps and often results in stronger performance.
-However, it also comes with challenges, such as costly updates—requiring fine-tuning whenever domain knowledge changes—and scalability issues, as embedding large or frequently changing knowledge bases can be computationally expensive.
+At inference time, no further retrieval or external knowledge calls are required: $\mathbf{y}=M\bigl{(}\mathbf{x};\Delta\theta\bigr{)}$.
+This paradigm enables fast inference by removing the need for additional retrieval steps and often delivers stronger performance.
+However, it also presents challenges, such as high update costs since fine-tuning is required when domain knowledge changes, and scalability concerns because embedding large or frequently changing knowledge bases demands significant computational resources.
 3.3 Modular Knowledge Adapters
@@
 $\phi=\textstyle\arg\min_{\phi}\sum_{(\mathbf{x_{s}},\mathbf{y_{s}})\in\mathcal{K}}\mathcal{L}\bigl{(}M(\mathbf{x_{s}};\theta,\phi),\mathbf{y_{s}}\bigr{)},$
-where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters. At inference time, the enhanced model generates outputs as:
-$\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}.$
+where $M(\mathbf{x_{s}};\theta,\phi)$ represents the base model’s generation function enhanced with the new adapter parameters.
+At inference time, the enhanced model generates outputs as: $\mathbf{y}=M\bigl{(}\mathbf{x};\theta,\phi\bigr{)}$.
 This paradigm offers a parameter-efficient method to adapt LLMs to specific domains without modifying the original model weights.
@@
 3.4 Prompt Optimization
-Unlike previous approaches, prompt optimization does not retrieve knowledge from external sources.
-Instead, it focuses on fully leveraging or guiding the LLM to utilize its internal, pre-existing knowledge.
+Prompt optimization refers to the practice of guiding LLMs to perform domain-specific tasks by crafting effective textual prompts.
+Unlike retrieval-based methods, it relies entirely on the model’s internal knowledge and does not require access to external knowledge bases or fine-tuning.
 The process can be formalized as:
-$\mathbf{y}=M\bigl{(}[\mathbf{p},\mathbf{x}];\theta\bigr{)},$
-where $\mathbf{p}$ represents a textual prompt containing implicit domain knowledge or specific instructions.
-Prompt optimization offers significant advantages, including eliminating dependency on external domain knowledge bases and avoiding training.
-However, it also presents challenges, as designing effective prompts can be both complex and time-consuming.
-Additionally, long prompts may reduce the available context window, potentially affecting model efficiency and performance.
+$\mathbf{p}^{*}=\textstyle\arg\min_{\mathbf{p}}\mathcal{L}\big{(}M([\mathbf{p},\mathbf{x}];\theta),\mathbf{y}^{*}\big{)},$
+where $\mathbf{p}$ is a prompt containing domain-relevant cues, $\mathbf{x}$ is the task input, and $\theta$ are the fixed parameters of the LLM.
+This paradigm offers practical advantages such as lightweight deployment, no training overhead, and adaptability across domains.
+However, it also faces key challenges: designing prompts that elicit accurate responses can be non-trivial, and long prompts may reduce the available input space due to context length limitations.
+Prompt-based approaches can be broadly categorized into manual prompting, prompt tuning, and prefix tuning.
+These differ in how prompts are constructed or optimized—ranging from discrete, static prompts to learnable embeddings—and have been widely adopted for low-resource domain adaptation.
 3.5 Comparison of the Four Paradigms
@@
 knowledge
-Table 2: Guidance on selecting knowledge injection paradigms based on training cost, inference speed, and limitations.
-Dynamic knowledge injection integrates external knowledge at runtime, offering flexibility and adaptability to new information without increasing training cost.
-However, it requires an effective retrieval module, and the inference speed depends heavily on retrieval performance, which can slow down the overall process.
-Static knowledge embedding embeds domain expertise during pretraining or fine-tuning, requiring large-scale domain-specific data and significant training resources, including GPUs and time.
-While it incurs no additional inference cost, its limitations lie in the potential risks of catastrophic forgetting and its inability to adapt to evolving information.
-Modular adapters serve as a middle ground, allowing plug-and-play components to enhance domain-specific capabilities with minimal training data.
-Only a small subset of parameters needs to be trained, which reduces training costs, and inference speed is largely unaffected.
-However, the quality of training data significantly impacts the performance of this method.
-Prompt Optimization, on the other hand, avoids retraining entirely by leveraging carefully crafted inputs.
-It has no impact on inference speed but relies on extensive human effort to find optimal prompts.
-This method is limited in its ability to utilize new knowledge and primarily activates pre-existing knowledge.
-We summarize these comparisons in Table 2 as a practical guide to help determine the most suitable method based on specific task requirements and scenarios.
+Table 2: Guidance on choosing injection methods by training cost, inference speed, and constraints.
+Dynamic knowledge injection introduces external knowledge at runtime, offering flexibility and adaptability without added training cost.
+However, it depends on efficient retrieval, and inference speed can suffer if retrieval performance is poor.
+Static knowledge embedding integrates domain expertise during pretraining or fine-tuning, requiring large-scale data and significant computational resources.
+It adds no inference cost but struggles to adapt to new information and is prone to catastrophic forgetting.
+Modular adapters offer a middle ground by enhancing domain capabilities through plug-and-play modules that require minimal training data.
+Only a small number of parameters are trained, reducing cost and preserving inference speed, though performance heavily depends on data quality.
+Prompt optimization avoids retraining by using well-crafted inputs.
+It maintains fast inference but relies on significant manual effort and is limited to activating existing knowledge rather than incorporating new information.
+We summarize these comparisons in Table 2 as a practical guide to help determine the suitable method based on specific requirements and scenarios.
 4 Applications
-4.1 Biomedicine
-The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts.
-These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data.
-For instance, PMC-LLaMA (Wu et al., 2023) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively.
-Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning.
-This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
-Beyond foundational models, integrating external tools and knowledge can further enhance performance.
-For example, GeneGPT (Jin et al., 2024) utilizes an LLM pretrained on code tasks to tackle GeneTuring tests by employing NCBI Web APIs.
-This approach combines in-context learning with an augmented decoding algorithm that identifies and executes API calls.
-Similarly, Med-PaLM (Singhal et al., 2023a) introduces vector prompts—representations that store and retrieve medical domain knowledge—to extend the capabilities of Flan-PaLM (Chung et al., 2024).
+4.1 Finance
+In the financial domain, LLM generally follow two main development paths: fine-tuning general-purpose models on financial tasks or training models from scratch using domain-specific corpora.
+For fine-tuning, PIXIU(Xie et al., 2023) adapts LLaMA using 136K financial instruction samples, equipping the model to handle diverse finance-related scenarios.
+Instruct-FinGPT(Zhang et al., 2023) focuses on sentiment classification by fine-tuning on 10K samples from two financial sentiment datasets.
+FinGPT (Yang et al., 2023) proposes an end-to-end framework for developing FinLLMs, efficiently fine-tuning LLaMA and ChatGLM with 50K samples via LoRA, significantly reducing computational costs.
+In contrast, scratch-trained FinLLMs aim for deep domain alignment from the ground up.
+BloombergGPT (Wu et al., 2023b) uses 5B Bloomberg-specific tokens (0.7% of its total corpus) to specialize in financial applications.
+XuanYuan 2.0 (Zhang and Yang, 2023) is the largest Chinese financial chatbot, trained on 366B tokens and fine-tuned on 13B.
+Fin-T5 (Lu et al., 2023) leverages a 300GB Chinese financial corpus using the T5 architecture, while SNFinLLM (Zhao et al., 2024a) enhances inference through real-time financial data injection.
+In summary, this field showcases a rich diversity in training strategies, from lightweight tuning to comprehensive, end-to-end development.
 Domain Model Paradigms Knowledge Source Link
-Biomedicine PMC-LLaMA (Wu et al., 2023) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
-Med-PaLM 2 (Singhal et al., 2023b) Static Knowledge Embedding MultiMed Link
-DALK  (Li and others, 2024)
-Dynamic Knowledge Injection
-Prompt Optimization
-MedQA, MedMCQA, MMLU, QA4MRE Link
-ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
-SA-MDKIF (Xu et al., 2024b) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
-MaLP (Zhang et al., 2024b) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
-BioMedLM  (Bolton et al., 2024) Static Knowledge Embedding PubMed,MedMCQA,MedQA,MMLU,BioASQ Link
-BiomedRAG (Li et al., 2024) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
-MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
-Finance FLANG  (Shah and others, 2022) Static Knowledge Embedding
+Finance FLANG  (Shah et al., 2022) Static Knowledge Embedding
 Financial PhraseBank,FiQA 2018 Task-1,
@@
 Link
-BloomBergGPT (Wu and others, 2023)
+BloomBergGPT (Wu et al., 2023b)
 Static Knowledge Embedding
@@
 Link
-Fin-LLaMA  (Konstantinidis and others, 2024) Static Knowledge Embedding fin-llama-dataset Link
+Fin-LLaMA  (Konstantinidis et al., 2024) Static Knowledge Embedding fin-llama-dataset Link
 SNFinLLM (Zhao et al., 2024a)
@@
 FinEval, FinanceIQ,qEQA,FinC,KQA,MRC,cMRC
+Fino1  (Qian et al., 2025) Static Knowledge Embedding
+FinQA, TATQA , DocMath (Simpshort and Compshort),
+DocFinQA , and BizBench-QA
+Link
+Biomedicine PMC-LLaMA (Wu et al., 2023a) Static Knowledge Embedding PMC-OA, MedC-I, PubMedQA, MedMCQA, USMLE Link
+Med-PaLM 2 (Singhal et al., 2023b) Static Knowledge Embedding MultiMed Link
+DALK  (Li et al., 2024a)
+Dynamic Knowledge Injection
+Prompt Optimization
+MedQA, MedMCQA, MMLU, QA4MRE Link
+ChronicCareGPT (Liu et al., 2024b) Prompt Optimization eRisk Link
+SA-MDKIF (Xu et al., 2024c) Modular Knowledge Adapters MedQuA,emrQA, PubMedQA, MedQA \
+MaLP (Zhang et al., 2024c) Modular Knowledge Adapters HealthCareMagic-100k, iCliniq Link
+BioMedLM  (Bolton et al., 2024) Static Knowledge Embedding PubMed,MedMCQA,MedQA,MMLU,BioASQ Link
+BiomedRAG (Li et al., 2024b) Dynamic Knowledge Injection CHEMPROT,DDI,ade-corpus-v2,MTsample,ADInt,UMLS Link
+MedINST (Han et al., 2024) Static Knowledge Embedding MedINST Link
+K-COMP (Cho and Lee, 2025) Dynamic Knowledge Injection MedCorp corpus \
+OntoTune (Liu et al., 2025c) Static Knowledge Embedding SemEval2018 Task 9 dataset Link
 Materials ChemCrow (Bran et al., 2023) Dynamic Knowledge Injection 18 expert-designed tools Link
@@
 Link
-ChemLLM (Zhang and others, 2024)
+ChemLLM (Zhang et al., 2024a)
 Static Knowledge Embedding
@@
 Link
-ScholarChemQA  (Chen and others, 2024)
+ScholarChemQA  (Chen et al., 2024)
 Static Knowledge Embedding
@@
 DARWIN 1.5 (Xie et al., 2024) Static Knowledge Embedding FAIR datasets Link
-ChemAgent (Tang and others, 2025)
+ChemAgent (Tang et al., 2025b)
 Dynamic Knowledge Injection
 Prompt Optimization
-Google API, Docker container, Internet,
-Hardware API document, Physical world hardware
+college chemistry textbooks:Quantum chemistry (quan), hemistry kinetics (matter)
+Quantum mechanics (chemmc), Physical chemistry (atkins)
 Link
+LLaMat  (Mishra et al., 2025) Static Knowledge Embedding MatBookQA,MaScQA,MatSciInstruct \
+OmniScience (Prabhakar et al., 2025) Static Knowledge Embedding daring-anteater dataset, s1K dataset \
 Mental Health MeChat (Qiu et al., 2023) Dynamic Knowledge Injection SMILECHAT, PsyQA Link
@@
 FPS (Liu et al., 2024e) Prompt Optimization Fake News Dataset, Big Five Personality Traits Link
-FUSE (Liu et al., 2024f) Prompt Optimization True News Dataset, Big Five Personality Traits \
+FUSE (Liu et al., 2025b) Prompt Optimization True News Dataset, Big Five Personality Traits \
 Table 3: Summary of the domain-specific knowledge injection studies. We categorize current work according to their research domain and knowledge injection method.
-4.2 Finance
-Fine-tuned financial LLMs have demonstrated significant advancements in adapting general-purpose models to domain-specific tasks through task-specific training.
-PIXIU (Xie et al., 2023) fine-tunes LLaMA on 136K instruction samples tailored to financial tasks, enabling the model to handle a wide range of domain-relevant scenarios.
-Instruct-FinGPT (Zhang et al., 2023) fine-tunes LLaMA on 10K instruction samples derived from two financial sentiment analysis datasets, focusing primarily on finance classification tasks.
-FinGPT (Yang et al., 2023) introduces a comprehensive end-to-end framework for training and deploying FinLLMs in the financial industry.
-Utilizing the LoRA technique, FinGPT fine-tunes open-source LLMs like LLaMA and ChatGLM with approximately 50K task-specific samples, achieving efficient fine-tuning without full model retraining.
-In contrast, scratch-trained financial LLMs aim to create models specifically designed for financial tasks from the ground up.
-BloombergGPT (Wu and others, 2023) leverages a subset of 5 billion tokens from Bloomberg-specific data, representing only 0.7% of its total training corpus, to tailor its model to financial applications.
-XuanYuan 2.0 (Zhang and Yang, 2023) combines 366 billion tokens for pre-training with an additional 13 billion tokens for fine-tuning, creating the largest Chinese financial chat model.
-Similarly, Fin-T5 (Lu et al., 2023) introduces a Chinese financial pre-training language model built on the T5 architecture, utilizing a 300GB financial corpus.
-Furthermore, SNFinLLM (Zhao et al., 2024a) dynamically incorporates real-time financial data during inference to enhance decision-making capabilities, demonstrating the value of domain-specific pretraining and adaptability in financial LLMs.
+4.2 Biomedicine
+The biomedicine domain benefits from a wealth of specialized corpora, such as PubMed (Dernoncourt and Lee, 2017) and MedQA (Jin et al., 2021), enabling the development of LLMs specifically trained on biomedical texts.
+These models often follow the static knowledge embedding approach, leveraging the domain-specific richness of biomedical data.
+For instance, PMC-LLaMA (Wu et al., 2023a) extends the LLaMA 7B model through further pretraining on 4.9 million PubMed Central articles curated from the S2ORC dataset (Lo et al., 2020), completing five epochs to embed biomedical knowledge effectively.
+Similarly, Med-PaLM 2 (Singhal et al., 2023b) builds on PaLM 2 via instruction fine-tuning.
+This fine-tuning incorporates a diverse mix of medical question-answering datasets, including MedQA, MedMCQA (Pal et al., 2022), and HealthSearchQA (Singhal et al., 2023a).
+Beyond foundational models, integrating external tools and knowledge can further enhance performance.
+GeneGPT(Jin et al., 2024) leverages a code-pretrained LLM to address GeneTuring tests by calling NCBI Web APIs, combining in-context learning with an augmented decoding algorithm capable of identifying and executing API requests.
+Med-PaLM(Singhal et al., 2023a) extends the capabilities of Flan-PaLM (Chung et al., 2024) through the use of vector prompts—dense representations designed to store and retrieve medical domain knowledge during inference.
+Overall, biomedical LLMs lead in combining static pretraining, instruction tuning, and tool integration, reflecting a shift toward hybrid reasoning in specialized AI.
 4.3 Materials
-In contrast to biomedicine, where significant efforts have been devoted to static knowledge embedding due to the availability of extensive corpora, research in materials and chemistry has primarily focused on utilizing task-related tools that align with the dynamic knowledge injection paradigm.
-For instance,  Xie et al.
-(2024) demonstrated how Darwin 1.5 leverages natural language inputs and a two-stage training strategy to achieve significant improvements in materials discovery and design tasks.
-Bran et al.
-(2023) introduced ChemCrow, a framework that augments LLMs with chemistry-expert-designed tools for downstream tasks such as organic synthesis and drug discovery.
-There are also studies on prompt optimization Tang and others (2025), demonstrating that better-designed planning prompts can effectively utilize the model’s internal knowledge to orchestrate complex tasks.
-This approach capitalizes on the planning and execution capabilities of multiple LLMs to achieve autonomy in chemical experimentation.
-More recently, there has been growing interest in exploring static knowledge embedding and modular knowledge adapters within the chemistry domain.
-For example, Chen and others (2024) curated a QA dataset to fine-tune pretrained models like BERT and LLMs such as Llama, aiming to enhance their performance in chemistry-related tasks.
-Similarly, Xie et al.
-(2024) introduced Darwin 1.5, an open-source large language model tailored for materials science.
+In contrast to the biomedical domain, the field of materials science and chemistry has largely focused on static knowledge embedding.
+Many models rely on domain-specific corpora to fine-tune general models for improved task performance.
+Darwin 1.5 (Xie et al., 2024) adopts a two-stage training strategy using natural language inputs to enhance performance in materials discovery.
+ScholarChemQA (Chen et al., 2024) constructs a chemistry QA dataset to fine-tune BERT and LLaMA, improving chemical reasoning.
+Recently, some efforts have begun to explore dynamic knowledge integration.
+ChemCrow (Bran et al., 2023) augments LLMs with chemistry tools for applications like synthesis and drug discovery.
+ChemAgent (Tang et al., 2025b) shows that well-designed planning prompts can guide models through complex execution tasks by leveraging internal reasoning.
+While still in early stages, the field is transitioning from static embedding toward interactive and tool-augmented reasoning, indicating strong potential for future developments
 4.4 Human-Centered Science
-The last domain we introduce is human-centered science, which encompasses a wide range of applications such as psychological counseling, financial forecasting, social behavior prediction, and legal reasoning.
-All these domains center on understanding and interacting with human needs, behaviors, and decision-making processes.
+Human-centered science focuses on understanding and assisting human behaviors, needs, and decisions.
+This interdisciplinary domain includes mental health, education, social behavior prediction, and legal reasoning—each benefiting from personalized and context-aware LLMs.
 In mental health, datasets like PsyQA (Sun et al., 2021) provide a foundation for training models in psychological counseling scenarios.
-For instance, SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations.
-Similarly, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, significantly enhancing its emotional support capabilities.
+SoulChat (Chen et al., 2023), a model fine-tuned on over 100,000 long-text counseling sessions using static knowledge embedding, is designed for empathic conversations.
+In contrast, MeChat (Qiu et al., 2023) employs dynamic knowledge injection to adapt to real-time inputs, enhancing its emotional support capabilities.
 These advancements demonstrate the potential of human-centered science in addressing complex, real-world challenges through personalized and context-aware solutions.
@@
 Personalized learning, for example, requires models to adapt to individual needs, providing tailored feedback and emotional support.
-EduChat (Dan et al., 2023) tackles this by leveraging educational theories from psychology and pedagogy through static knowledge embedding, enabling tasks like open Q&A, composition correction, and emotional support.
-Similarly, QiaoBan (Weixiang et al., 2023) focuses on child-centered education by using prompt optimization to adapt the model’s behavior based on child psychology and emotional well-being, catering specifically to young learners.
-Domain-specific education and interactive teaching have also seen advancements through LLMs.
+EduChat (Dan et al., 2023) applies psychological and pedagogical theories via static knowledge embedding to support tasks like Q&A, writing feedback, and emotional guidance.
+Similarly, QiaoBan (Weixiang et al., 2023) uses prompt optimization to tailor model behavior to children’s psychological and emotional needs.
+Domain-specific education and interactive teaching have also seen advancements.
 CyberQ (Agrawal et al., 2024) blends static knowledge embedding and dynamic knowledge injection via AISecKG (Agrawal, 2023), generating Q&A based on cybersecurity best practices.
@@
 For social sciences, models like SocialLLM (Jiang and Ferrara, 2023) combine static knowledge embedding and dynamic knowledge injection to analyze human behavior in social networks.
-Adapters facilitate large-scale data integration while prompt optimization guides the model to focus on specific social behavior patterns.
-Models like FPS (Liu et al., 2024e) and FUSE (Liu et al., 2024f) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact.
-A summary of the mainstream models and their information is provided in Table 3. More models across various domains can be found at: Survey-official-repo.
+Models like SSF (Wang et al., 2024a) FPS (Liu et al., 2024e) and FUSE (Liu et al., 2025b) use prompt optimization to simulate the spread and evolution of fake news in social networks, helping understand misinformation’s impact.
+In addition, Liu et al.
+(2025a) adopts a multi-agent approach to synthesize data, which is then used as knowledge to enhance language models for detecting fake news.
+A summary of the mainstream models and their information is provided in Table 3.
+More models across various domains can be found at: official-repo.
 5 Tools, Resources, and Analysis
 5.1 Knowledge Injection Framework
-In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024c) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
+In this section, we provide a detailed introduction to four open-source frameworks categorized under different knowledge injection methods to facilitate understanding and application: KnowGPT (Zhang et al., 2024d) for Dynamic Knowledge Injection, StructTuning (Liu et al., 2024d) for Static Knowledge Embedding, K-Adapter (Wang et al., 2021) for Modular Knowledge Adapters, and SelfLift (Cheng et al., 2024) for Prompt Optimization.
 KnowGPT dynamically combines knowledge graphs with prompt optimization by leveraging reinforcement learning to extract highly relevant subgraphs from the knowledge graph.
@@
 Finally, SelfLift iteratively employs a retrieval-augmented generator to create an unbounded memory pool and uses a memory selector to choose one output as memory for the subsequent generation round.
-This is an excellent demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
-5.2 Datasets and Benchmarks
-We summarize commonly used datasets or benchmarks for domain-specific LLM study in Table 3, observing significant variation in dataset richness across domains.
-Biomedicine boasts numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis and others (2012), that support tasks such as question answering and clinical summarization.
+This is a good demonstration of prompt optimization, where the model’s outputs are dynamically refined and reused to enhance its overall performance and coherence in subsequent tasks.
+5.2 Knowledge Source
+We summarize commonly used knowledge sources for domain-specific LLMs in Table 3, referring to datasets that provide the external knowledge used in various injection methods—including training corpora for static embedding or adapter tuning, and retrieval or prompt design resources for dynamic knowledge injection.
+Biomedicine includes numerous high-quality datasets, such as PubMed, PubMedQA (Jin et al., 2019), and BioASQ Tsatsaronis et al.
+(2012), which support tasks such as question answering and clinical summarization.
 In contrast, materials and chemistry have more limited resources, and datasets like USPTO and Enzymes focus on chemical reactions.
@@
 This diversity underscores the effort to tailor LLMs to specialized fields while emphasizing the need for broader curation of benchmarks in underrepresented domains.
-5.3 Performance Comparison of Domain-specific LLM and General-domain LLM
-Model Domain Type Size MedQA PubMedQA MedMCQA
-Med-Gemini Specific 91.1 - -
+5.3 Performance Comparison of 4 Paradigms
+Model Category MedQA PubMedQA MedMCQA
+GPT-4 (Medprompt) Prompt Optimization 90.2 82.0 79.1
 GPT-4 General 90.2 80.4 73.7
-Med-PaLM 2 Specific 85.4 81.8 72.3
-PMC-LLaMA Specific 13B 56.3 77.9 56.0
-BioMedLM Specific 2.7B 50.3 74.4
-Llama 2 General 70B 43.7 74.3 35.0
-Galactica General 120B 44.4 77.6 52.9
-Table 4: Performance comparison of domain-specific and general-domain model performance on medical benchmarks.
-Since there are also general-domain strong LLMs, it is necessary to discuss the comparison between domain-specific LLMs and general ones to determine if a specific knowledge injection process is essential.
-Here, we take the biomedical domain as an example due to significant research efforts in this field, as shown in Table 4.
-The results are collected from corresponding papers or paperswithcode.com.
-First, we can observe that closed-source LLMs are currently the most effective models, while the performance gap between general-domain and domain-specific LLMs is relatively narrow.
-For example, both GPT-4 and Med-Gemini (Saab et al., 2024) achieve outstanding performance, with scores higher than 90 on the MedQA dataset.
-However, due to the lack of transparency in closed-source LLMs, open-source LLM efforts should not be overlooked.
-In this field, domain-specific LLMs often outperform general-domain models.
-For instance, PMC LLaMA-13B outperforms LLaMA2-70B by more than 10 points on the MedQA dataset.
-This demonstrates the value of domain-specific LLMs in achieving superior performance on specialized tasks.
-While general-domain models can deliver strong results, incorporating domain-specific knowledge allows for significant improvements, particularly in open-source initiatives.
-This underscores the importance of investing in domain-specific LLMs to address the unique challenges of specialized fields.
+Med-PaLM 2 Static Knowledge 85.4 81.8 72.3
+Flan-PaLM (3-shot) Dynamic Knowledge 67.6 79.0 57.6
+PMC-LLaMA Static Knowledge 56.3 77.9 56.0
+BioMedLM Static Knowledge 50.3 74.4
+LLaMA (MedAdapter) Knowledge Adapters 37.4 63.6 32.0
+Table 4: Model performance across four knowledge paradigms on medical benchmarks.
+To compare knowledge injection paradigms in a practical setting, we focus on the biomedical domain due to its popularity and the availability of benchmarks such as MedQA, PubMedQA, and MedMCQA, as shown in Table 4.
+Although the models differ in architecture, we align backbones when possible.
+For example, both PMC-LLaMA and MedAdapter use LLaMA-13B.
+SOTA models like GPT-4 are closed-source, making prompt optimization the only feasible adaptation strategy.
+Despite no domain-specific training, GPT-4 with Medprompt achieves strong performance, showing the effectiveness of prompt methods for closed models.
+Among open models, MedAdapter underperforms compared to PMC-LLaMA, suggesting that full fine-tuning may outperform adapter-based methods for some tasks.
+Performance differences across paradigms also highlight the importance of pretraining corpus and task alignment, particularly in static injection approaches.
+Furthermore, in Appendix A, we also compare knowledge injection paradigms in the finance domain and obtain similar conclusions to those in the medical domain.
 6 Challenges and Opportunities
-6.1 Integrated Knowledge Consistency
+Integrated Knowledge Consistency.
 Knowledge injection allows LLMs to incorporate and integrate different domain-specific knowledge.
-However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024a).
-For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise, resulting in unreliable decisions and undermining the system’s trustworthiness.
+However, retrieved knowledge may conflict with the model’s pre-trained representations or other retrieved facts, leading to inconsistencies in outputs (Xu et al., 2024b).
+For example, in healthcare or legal analysis, conflicting treatment protocols or contradictory legal precedents could arise Dayton (2012), resulting in unreliable decisions and undermining the system’s trustworthiness.
 To address this, future research must focus on detecting inconsistencies, resolving conflicts, and maintaining consistency in integrated knowledge.
 Conflicts can be addressed by prioritizing reliable sources, applying domain-specific rules, or using ensemble techniques to balance multiple perspectives.
-Alignment algorithms and validation modules can further ensure that retrieved knowledge aligns with the model’s reasoning processes and is reliable before influencing outputs.
-These efforts are essential to enhance the reliability and applicability of knowledge-enhanced LLMs in complex, high-stakes domains.
-6.2 Cross-Domain Knowledge Transfer
-Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields.
-While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns.
-For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks.
-Overcoming these challenges necessitates advancements in modular knowledge representation and transfer learning techniques.
+Alignment and validation modules help ensure retrieved knowledge fits the model’s reasoning.
+Cross-Domain Knowledge Transfer.
+Cross-domain knowledge transfer involves equipping LLMs with the ability to generalize knowledge across diverse and distinct fields Li et al.
+(2025b).
+While this significantly expands their applicability, it also introduces challenges due to the complexity and diversity of domain-specific terminologies, ontologies, and reasoning patterns Montero et al.
+(2004).
+For example, transferring knowledge from chemistry to healthcare might require reconciling differing data structures and reasoning frameworks (Schroeder et al., 2018).
+Overcoming these challenges requires advancements in modular knowledge representation and transfer learning techniques.
 Future efforts could explore hybrid approaches that blend static embeddings with dynamic retrieval, enabling LLMs to adapt knowledge flexibly across domains without compromising depth.
-Additionally, creating standardized cross-domain benchmarks and datasets could facilitate systematic evaluation and foster innovation in multi-domain knowledge transfer methodologies.
+Additionally, standardized cross-domain benchmarks can enable consistent evaluation and drive innovation in knowledge transfer methods.
+We provide more discussions in Appendix B.
 7 Conclusion
@@
 We also summarize applications across biomedicine, chemistry, and computational social science domains.
-By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires further exploration of knowledge-enhanced LLMs for domain-specific challenges.
+By highlighting standard datasets, benchmarks, challenges, and future opportunities, we aim to provide a valuable resource that inspires exploration of knowledge-enhanced LLMs for domain-specific challenges.
+Limitation
+Despite providing a comprehensive review of current methods and applications for domain-specific knowledge injection in LLMs, this survey has certain limitations.
+First, while we strive to cover several key domains such as finance, biomedicine, and materials science, some less-studied or emerging areas (for example, low-resource languages, cross-cultural education, and niche disciplines) receive relatively limited attention.
+Second, our focus is primarily on summarizing methodological principles and representative models from existing literature.
+Due to substantial variation in model architectures, application domains, training data, and evaluation protocols, we were only able to conduct targeted comparisons under controlled conditions within the biomedical domain, using commonly adopted datasets.
+A more systematic and broad-based empirical evaluation across methods remains an important direction for future work.
+Nevertheless, we hope this survey serves as a useful reference and provides a clear roadmap for ongoing research in knowledge-enhanced LLMs.
+Acknowledgement
+We would like to thank the anonymous reviewers for their constructive comments.
+The work was supported by Mohamed bin Zayed University of Artificial Intelligence (MBZUAI) through grant award 8481000078.
+Appendix A Performance Comparison of 4 Paradigms
+Model Category FPB FiQA-SA TFNS
+GPT-4 General 83.3 63.0 80.8
+GPT-4(finetune) Static Knowledge 87.8 88.7 88.3
+FinGPT Knowledge Adapters 88.2 87.4 90.3
+FinBERT Static Knowledge 88.0 59.6 73.3
+BloombergGPT Static Knowledge 51.1 75.1 -
+Llama2-7B General 39.0 80.0 29.6
+FLANG Static Knowledge 91.9 3.4 -
+Table 5: Performance comparison of representative models under the four knowledge injection paradigms on financial benchmarks.
+To systematically compare the effectiveness of different knowledge injection paradigms in practical settings, we focus on two representative domains: biomedicine and finance.
+In the biomedical domain, which is widely studied and rich in benchmark datasets (e.g., MedQA, PubMedQA, and MedMCQA), we evaluate the performance of various models (see Table 4).
+Although the models differ in architecture, we align their backbones where possible—for example, both PMC-LLaMA and MedAdapter use LLaMA-13B.
+For closed-source models like GPT-4, prompt engineering is the only feasible adaptation strategy.
+Nevertheless, GPT-4 combined with Medprompt achieves strong performance, demonstrating the effectiveness of prompt-based knowledge injection.
+Among open models, MedAdapter underperforms compared to fully fine-tuned models such as PMC-LLaMA, suggesting that full fine-tuning may be more effective than adapter-based methods for certain tasks.
+Models with static knowledge (e.g., MedBERT) show substantial variance across tasks, underscoring the importance of alignment between pretraining corpora and downstream objectives.
+In Table 5, we extend this comparison to the financial domain, evaluating models on benchmarks such as FPBMalo et al.
+(2014), FiQA-SAMaia et al.
+(2018), and TFNSEl-Haj et al.
+(2020).
+The findings closely mirror those in the biomedical domain.
+Finetuned GPT-4 consistently outperforms others, confirming the value of injecting domain-specific knowledge into general-purpose LLMs.
+Static knowledge models like FinBERT and FLANG perform well on certain tasks but show significant variability, again emphasizing the crucial role of corpus-task alignment.
+FinGPT, which adopts lightweight adapter-based knowledge injection, achieves competitive performance while maintaining adaptability.
+In contrast, LLaMA2-7B lags behind across most tasks, reinforcing the necessity of targeted knowledge injection for domain-intensive applications.
+The consistency of observations across both domains suggests that the effectiveness of knowledge injection depends on a careful balance of architectural design, adaptation strategy, and corpus alignment to support complex, high-stakes tasks.
+Appendix B Detailed Discussions on Challenges & Opportunities
+B.1 Integrated Knowledge Consistency
+While knowledge injection empowers LLMs to reason with external facts, it introduces a crucial consistency problem: injected knowledge may contradict either the model’s internal representations or other pieces of retrieved information Xu et al.
+(2024b).
+In high-stakes domains such as healthcare and law, even minor inconsistencies can lead to significant consequences—for instance, conflicting drug dosages from different clinical guidelines (Dayton, 2012; Zhao et al., 2025), or divergent legal interpretations across jurisdictions (Guha et al., 2023).
+Recent research proposes techniques such as post-retrieval contradiction detection (Xu et al., 2024a), and confidence-aware re-ranking (Ren et al., 2025) to address these issues.
+Some frameworks, like MedRAG(Zhao et al., 2025), apply weighted retrieval and ensemble voting to prioritize reliable sources.
+Others explore neural symbolic consistency checking, where injected knowledge is aligned to pre-defined ontologies or verified using structured reasoning paths (Ciatto et al., 2024).
+Another growing area involves alignment-aware reranking, where retrieved documents are filtered based on their alignment with the LLM’s intermediate beliefs (Jin et al., 2025).
+Future directions may include interactive consistency resolution (e.g., user-in-the-loop conflict selection), as well as integrating factual calibration modules (Dong et al., 2022) that explicitly monitor factuality during decoding.
+These methods collectively aim to make knowledge-enhanced LLMs more robust, explainable, and reliable in dynamic or sensitive environments.
+B.2 Cross-Domain Knowledge Transfer
+Cross-domain transfer is a central challenge in building generalized yet specialized LLMs.
+As LLMs are exposed to knowledge from diverse domains, they must navigate incompatible ontologies, varied domain languages, and distinct reasoning structures Li et al.
+(2025b).
+For instance, transferring concepts from chemistry to healthcare involves not only bridging terminology gaps but also adapting to different causal assumptions and data formats (Schroeder et al., 2018).
+Several strategies have been proposed to manage this complexity.
+Adapter-based modularization He et al.
+(2021) allows domain-specific components to be trained separately and selectively activated.
+Meta-learning approaches(Hou et al., 2022) help models rapidly adapt to new domains with minimal supervision.
+Additionally, continual pretraining on mixed-domain corpora (Jin et al., 2022) offers a scalable method to improve robustness without catastrophic forgetting.
+Standardized datasets such as CrossNER(Liu et al., 2021) for multilingual named entity recognition, MultiLexSum(Shen et al., 2022) for cross-domain summarization, and MEDIQA-QA (Yadav et al., 2021) for biomedical QA serve as valuable testbeds for cross-domain evaluation.
+Future work may explore retrieval-augmented transfer, where dynamic selection of domain-relevant knowledge supports adaptive reasoning, or domain-invariant embedding learning, enabling LLMs to generalize across tasks without explicit supervision.
+B.3 Scalability and Efficiency of Knowledge Integration
+As LLMs are increasingly augmented with large-scale external knowledge such as entire knowledge graphs, document corpora, or real-time retrieval APIs, the computational and memory cost of incorporating this knowledge becomes a bottleneck.
+Efficient integration remains a key challenge, especially when operating under low-resource or latency-constrained settings.
+Techniques such as sparse retrieval (Lee et al., 2019), memory compression (Zhong et al., 2024), and caching strategies have been proposed to reduce overhead.
+Modular architectures (e.g., adapters or plug-in modules) also allow partial activation of knowledge, improving scalability.
+Future research could explore task-aware pruning of external knowledge, knowledge distillation from retrieval-based pipelines into compact models, and efficient routing mechanisms to select only relevant knowledge for each input.
+B.4 Evaluation and Hallucination Detection
+Evaluating knowledge-enhanced LLMs remains difficult due to the lack of standardized benchmarks and automatic metrics for factual consistency, coverage, and reasoning depth.
+Moreover, LLMs often hallucinate facts even when augmented with accurate knowledge (Ji et al., 2023), making it hard to trust outputs in high-stakes tasks.
+Recent work explores metrics like FactScore (Min et al., 2023), entailment-based verification (Patwa et al., 2022), and human-in-the-loop evaluation schemes.
+However, few of these methods scale across domains or languages.
+There is a growing need for task-specific, fine-grained evaluation metrics that capture whether the model used the retrieved knowledge effectively and truthfully.
+Additionally, incorporating hallucination detection as an internal module, such as through consistency checks between generation and source knowledge, may help reduce risk and improve interoperability.